OpenStudy (anonymous):

Simplify (2x^5 y^5 z^4)^5

OpenStudy (anonymous):

2^5=32 (x^5)^5=x^(5*5)=x^25 (y^5)^5=y^25 (z^4)^5=z^20 (2x^5 y^5 z^4)^5 =32x^25 y^25 z^20

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

you are welcome :D